Suppose there is an x-linked recessive disease linked to baldness. In a large randomly mating population of 1600 females, 1 is bald from the genetic disorder. What is the allelic frequency of males with hair.
Answer #1

Answer:

Frequency of bald females =q^2=1/1600= 0.000625

Frequency of baldness allele=q= sqrt of 0.000625= 0.025

The normal allele = p= 1-0.025= 0.975

The frequency of allele = The frequency of phenotype in males (as males have only one chromosome)

Therefore the frequency of males with hair = 0.975